OpenStudy (anonymous):

make up an algebra problem involving two values which must be solved. then solve it, showing all steps and the solution.

OpenStudy (turingtest):

The classic: I have 4 coins in my hand, all quarters and dimes they total a value of 85cents how many of each coin do I have?

OpenStudy (anonymous):

The length of a rectangle is 7cm more than twice the width of the rectangle. If the perimeter of the rectangle is 38cm find the length and width of the rectangle. Let the length of the rectangle be x and width be y. Then x=2y+7 2x+2y=38--eqn1 x-2y=7--eqn2 eqn1+eqn2=>3x=45 x=15 2y=x-7 2y=15-7 =8 y=4 The length of the rectangle is 15cm and width of the rectangle is 4cm.